OpenStudy (nwolf):

What is the slope of (-3,2) M=???

OpenStudy (anonymous):

y=x+b 2=-3+b b=5 I'm not sure.. Wait till someone else confirms if I'm right or not :S

OpenStudy (anonymous):

y=mx+b is the equation for the slope of a line. You will need two points to find the slope between those two points. A single point, (-3,2) will not give you a slope